# Politecnico di Torino, DIMEAS, Corso Duca degli Abruzzi, 24 10129 Torino, Italy

### v1

@@ Abstract
In chaotic dynamical systems, extreme events manifest in time series as unpredictable large-amplitude peaks. Although deterministic, extreme events appear seemingly randomly, which makes their forecasting difficult. By learning the dynamics from observables (data), reservoir computers can time-accurately predict extreme events and chaotic dynamics, but they may require many degrees of freedom (large reservoirs). In this paper, by exploiting quantum-computer ansätze and entanglement, we design reservoir computers with compact reservoirs and accurate prediction capabilities. First, we propose the recurrence-free quantum reservoir computer (RF-QRC) architecture. By developing ad-hoc quantum feature maps and removing recurrent connections, the RF-QRC has quantum circuits with small depths. This allows the RF-QRC to scale well with higher-dimensional chaotic systems, which makes it suitable for hardware implementation. Second, we forecast the temporal chaotic dynamics and their long-term statistics of low- and higher-dimensional dynamical systems. We find that RF-QRC requires smaller reservoirs than classical reservoir computers. Third, we apply the RF-QRC to the time prediction of extreme events in a model of a turbulent shear flow with turbulent bursts. We find that the RF-QRC has a longer predictability than the classical reservoir computer. The results and analyses indicate that quantum-computer ansätze offer nonlinear expressivity and computational scalability, which are useful for forecasting chaotic dynamics and extreme events. This work opens new opportunities for using quantum machine learning on near-term quantum computers.
@@ I Introduction:
Data-driven prediction of chaotic dynamical systems has gained significant interest in the last decade Cheng et al. (2015); Ghadami and Epureanu (2022). A data-driven model learns the solution of a dynamical system from data with the goal of predicting its temporal evolution. Weather and climate predictions Holmstrom et al. (2016); Rolnick et al. (2022), financial time-series forecasting Sezer et al. (2020), thermoacoustics Huhn and Magri (2020a), turbulence Srinivasan et al. (2019); Doan et al. (2019a), among many others, are examples of chaotic dynamical systems. The use of data-driven methods for predicting nonlinear systems has been motivated mainly by two reasons: (a) the availability of large amounts of data for these systems, and (b) the difficulty of predicting temporal dynamics due to chaos. In chaotic systems, a small change in the initial conditions can drastically change the dynamical system’s solution.
@@
Because chaotic dynamics is a sequential-data problem, recurrent neural networks (RNNs) are a natural choice for chaotic forecasting. RNNs are machine learning models that introduce recurrence in conventional neural network architectures. RNNs are employed in time-series forecasting, modeling nonlinear dynamical systems and chaos, and extreme event predictions Vlachas et al. (2020); Srinivasan et al. (2019); Racca and Magri (2022a); Doan et al. (2019b); Margazoglou and Magri (2023); Racca and Magri (2021). On the one hand, the computational capabilities of RNNs are excellent. On the other hand, RNNs are difficult to train as they require backpropagation through time at each time step Werbos (1990). Because RNNs are designed to learn correlations using internal hidden states and long-lasting time dependencies, the training through backpropagation can be difficult for long time-series forecasting Vlachas et al. (2020).
@@
Reservoir Computing (RC) bypasses the problem of backpropagation by introducing a static recurrence—the reservoir—and training with ridge regression. This avoids backpropagation at each time step. Reservoir computing is a unified computing framework that was introduced as ’Echo State Networks’ Jaeger (2001) and ’Liquid State Machines’ Maass et al. (2002). In this study, we work with Echo State Networks, which are a type of reservoir computing approach because of their potential in forecasting chaotic dynamics Doan et al. (2019b); Vlachas et al. (2020); Boedecker et al. (2012), extreme events predictions Doan et al. (2019c); Racca and Magri (2022b); noa (a) and the stability properties Margazoglou and Magri (2023); Huhn and Magri (2020a). (For brevity, we will refer to echo state networks as reservoir computing in this work). Despite their excellent forecasting abilities, the prediction capabilities of reservoir computers are sensitive to the choice of hyperparameters and the reservoir size, which is constrained by the computational resources and memory of current classical computers.
@@
Improving the performance of reservoir computing typically requires higher computational resources. Quantum computers hold the potential to provide exponential speed-up over classical computers for certain computational tasks Feynman (2018). This speed-up is also known as ’quantum advantage’. There are already algorithms that have theoretically proven this concept. For example, the prime factorization problem, which is the basis of most modern encryption and is NP-hard classically, can be solved in polynomial time using quantum computing principles Shor (1994). The second example is the matrix inversion problem for solving linear systems of equations, which is a key step in many numerical solvers. Proposed quantum algorithms for matrix inversion have the potential to provide quadratic speed-up over conventional classical methods Harrow et al. (2009); Bravo-Prieto et al. (2020). The underlying phenomena that enable the quantum advantage are entanglement, superposition, and interference effects in quantum computers. Recently, quantum computing has been applied in the area of machine learning. The current quantum hardware, however, is limited by the maximum number of qubits in the order of $\sim$ $10^{2}-10^{3}$ with environmental noise and decoherence. In the current Noisy Intermediate Scale Quantum (NISQ)-era Bharti et al. (2022), hybrid quantum-classical methods are one of the leading candidates with prospects of achieving a quantum advantage by employing methods such as Variational Quantum Algorithms and Quantum Circuit Learning Cerezo et al. (2021); Mitarai et al. (2018). The hybrid quantum-classical methods that can be used for the prediction of chaotic dynamics and extreme events are inspired by a branch of classical machine learning known as Recurrent Neural Networks (RNNs).
@@
Quantum Reservoir Computing (QRC) is a type of Quantum Machine Learning model that combines both frameworks, including Quantum Computing and Reservoir Computing. The idea of QRC is to use quantum dynamics to enhance the reservoir implementation Fujii and Nakajima (2017, 2021); Chen et al. (2022). Some numerical implementations of QRC have shown that the quantum systems of 5-7 qubits have comparable computational capabilities to the classical reservoir sizes of 100-500 Fujii and Nakajima (2017). Previous works have proposed hybrid QRC architectures for time series predictions and chaos modeling Pfeffer et al. (2022); Suzuki et al. (2022); Fry et al. (2023); Wudarski et al. (2023). In hybrid QRC, the input data is encoded in form of Bloch-sphere rotation angles in individual qubits along with reservoir states, which are the probability amplitudes. The resulting quantum states undergo a unitary operation at each time step Pfeffer et al. (2022). In another approach, a reservoir state is associated with the density operator of the encoded quantum state, and reservoir states correspond to measured expectation values for each qubit Fry et al. (2023); Dudas et al. (2023). Although it has been conjectured that quantum reservoir computers are potential candidates for providing a quantum advantage in the near-term NISQ era, only a few studies have been conducted on analysing higher-dimensional chaotic systems Tran and Nakajima (2020), and long-term statistical predictions using QRC. The reservoir size in classical reservoir computers is often a bottleneck in higher-dimensional systems. Reservoir size is associated with the size of the matrix that needs to be inverted in the training phase and is limited by the classical computer’s memory Vlachas et al. (2020), which also limits the predictability of classical reservoir computing methods. Consequently, using QRC to predict high-dimensional nonlinear dynamical systems that require a large reservoir size is key to realizing a quantum advantage of QRC in predicting chaos.
@@
In this work, we build up on the concept of hybrid QRC that encodes reservoir states as probability amplitudes of the quantum system Pfeffer et al. (2022). We use gate-based quantum computing to implement this architecture and use it for high-dimensional chaotic time-series forecasting and extreme-event predictions. More specifically, we investigate the 3-dimensional Lorenz-63, 10 Lorenz-96 systems Lorenz (1995) and the MFE (Moehlis, Faisst, and Eckhardt) Moehlis et al. (2004) model for time-accurate, statistical predictions, and extreme events forecasting. Extreme events are sudden and unmitigated changes in the observables of chaotic flows, while long-term statistical prediction is also an important metric in forecasting chaos.
@@
We benchmark our classical and quantum reservoir networks by studying them for low-order models and then extend our analysis to incorporate higher-dimensional dynamics. Both classical and quantum networks are used for extreme event predictions in a reduced order model of a shear flow between infinite plates subjected to a sinusoidal body forcing, also referred to as ’Moehlis, Faisst, and Eckhardt (MFE)’ Moehlis et al. (2004). We assess the performance using multiple performance metrics: first, by considering short-term time series prediction capabilities that are quantified using the Predictability Horizon (PH) and Valid Prediction Time (VPT) Vlachas et al. (2020); Boffetta et al. (2002), and, second, by comparison of long-term statistics that are quantified using statisical measures such as Probability Density Function (PDF) and F-Score Margazoglou and Magri (2023); Racca and Magri (2021).
@@
This paper is structured as follows. In Sec. II, we present the classical and proposed quantum reservoir architectures and their comparison. Different quantum reservoir architectures refer to the different ansätze chosen for forming a reservoir. Second, we apply both classical and quantum reservoir architectures on 3-dimensional Lorenz-63, 10-, and 20-dimensional Lorenz-96 systems. We increase the dimensionality and complexity of the dynamical systems progressively for the analysis. Third, we compare the results of the best quantum feature map with the best classical reservoir applied to the MFE model to analyse the extreme-event prediction capabilities of both architectures. The numerical models of different dynamical systems and the results for the reservoir predictions are presented in Sec. IV. Finally, in Sec. V, we conclude our findings by highlighting the potential of QRC in chaotic time series forecasting along with its associated challenges and the direction of future works.
@@ II.1 Classical Reservoir Computing (CRC)
Unlike other RNNs, reservoir computing is a type of RNN framework that does not require backpropagation through time, which represents a major computational advantage Jaeger (2001); Racca and Magri (2021). As a result, the training cost is substantially reduced and the objective function is minimized through a simple linear ridge regression.
Figure 1: Schematic representation of a classical reservoir computer Jaeger (2001). The input data $\boldsymbol{u}_{in}$ is mapped to the reservoir matrix via $\boldsymbol{W}_{in}$. The reservoir neuron connections governed by $\boldsymbol{W}$ matrix allow the flow of information between neurons. The linear readout layer using the trained $\boldsymbol{W}_{out}$ matrix is used to make output predictions $\boldsymbol{u}_{p}$.
Figure 2: Schematic Representation of reservoir computers. On the left-hand side, is the open loop training phase. On the right side, is the closed loop prediction phase.
@@
Let $\{\boldsymbol{\hat{u}}(t_{0}),\boldsymbol{\hat{u}}(t_{1}),\ldots,\boldsymbol{\hat{u}}(t_{N_{tr}})\}$ be a training data set of a dynamical state $\hat{\boldsymbol{u}}\,\epsilon\,\,\mathbb{R}^{{N}_{u}}$ , which is known at $N_{tr}+1$ steps in time. Generally, it is recommended to choose a set of variables that is (re-)scaled by the range component-wise, which is indicated by ( $\hat{}$ ) on the training data set Lukoševičius (2012). In reservoir computing, these states are mapped onto higher-dimensional reservoir states of an $N_{r}$-dimensional vector space. For a randomly chosen initial reservoir state $\boldsymbol{r}(t_{0})$, we recursively compute a set of $N_{tr}$ reservoir vectors $\boldsymbol{r}(t_{1}),\ldots,\boldsymbol{r}(t_{N_{tr}})$ as
$\boldsymbol{\hat{r}}(t_{i+1})$ $=\tanh\,(\boldsymbol{W}_{in}\boldsymbol{\hat{u}}_{in}(t_{i})+\boldsymbol{W}\boldsymbol{r}(t_{i})),$ (1)
$\boldsymbol{r}(t_{i+1})$ $=(1-\epsilon)\,\boldsymbol{r}(t_{i})+\epsilon\,\boldsymbol{\hat{r}}(t_{i+1}).$ (2)
@@
Here, the state $\hat{\boldsymbol{r}}\,\epsilon\,\,\mathbb{R}^{{N}_{r}}$ represent a state-vector containing reservoir activation states, $\boldsymbol{W}_{in}\,\epsilon\,\,\mathbb{R}^{{N}_{r}\,\times{N}_{u}}$ is the input matrix and $\boldsymbol{W}\,\epsilon\,\,\mathbb{R}^{{N}_{r}\,\times{N}_{r}}$ is the reservoir weight matrix Jaeger (2001). The $\tanh$ function is applied component-wise to provide a nonlinear activation. The matrices $\boldsymbol{W}_{in}$ and $\boldsymbol{W}$ are pseudo-randomly generated and constant throughout the training and prediction phase. The elements of $\boldsymbol{W}_{in}$ are sampled from a uniform distribution in $[-{\sigma_{in}},{\sigma_{in}}]$, where ${\sigma_{in}}$ is the input-scaling. The reservoir weight matrix $\boldsymbol{W}$ is an Erdös-Rényi matrix (Erdös-Rényi is a model for generating random graphs or the evolution of a random network in the field of graph theory), with average connectivity between each neuron varied by input reservoir density $D$. The spectral radius of the reservoir state matrix $\rho$ is given by its maximal eigenvalue. The matrix $\boldsymbol{W}$ is rescaled such that $\rho\leq 1$ Lukoševičius (2012); Jaeger (2001). Effectively, this guarantees a rapid decay of temporal correlations between successive reservoir states and helps to avoid overfitting. Whereas Eq. (1) is the nonlinear activation step, Eq. (2) combines the linear memory of the reservoir with the nonlinear activation, as parametrized by the leaking rate $\epsilon$. This additional post-processing step is known as memory non-linearity tradeoff Lukoševičius (2012).
Reservoir computing networks can be run in either open-loop or closed-loop configurations (Fig. 2). In the open loop, we use the input data at each time step and compute the corresponding reservoir dynamics $\boldsymbol{r}(t_{i})$ according to Eqs. (1) and (2). Because of the initialization process with a randomly chosen reservoir state, it is necessary to discard a small number $N_{w}$ of the initial transients to satisfy the Echo State Property (ESP) Jaeger (2001), which leads to dynamics that are less sensitive to the initial random choice. The process of discarding the initial transients is also called the washout phase. After the washout interval, reservoir dynamics at each time step $\boldsymbol{r}(t_{i})$ are collected to form a reservoir state matrix $\boldsymbol{R}\,\,\epsilon\,\,\mathbb{R}^{{N}_{r}\,\times(N_{tr}-N_{W})}$. Reservoir computing in the closed loop prediction mode requires an output matrix $\boldsymbol{{W}}_{out}$ for forecasting the dynamics of the learned chaotic system. The training of the output matrix $\boldsymbol{{W}}_{out}$ involves minimizing the mean square error between input and output data over the training data set, consisting of $N_{tr}$ number of training steps. The simplicity of the reservoir network allows us to achieve this by solving a linear ridge regression problem
$(\boldsymbol{R}\boldsymbol{R}^{T}+\beta\boldsymbol{I})\,\boldsymbol{W}_{out}=\boldsymbol{R}\,\boldsymbol{U}^{{\textnormal{T}}}_{d},$ (3)
@@
where $\beta$ is a user-defined Tikhonov regularization parameter to prevent overfitting, $\boldsymbol{I}$ is the identity matrix, and $\boldsymbol{{U}}^{{\textnormal{T}}}_{d}\,\,\epsilon\,\,\mathbb{R}^{{N}_{u}\,\times{N}_{tr}}$ is the horizontal concatenation of the output data. Once the output matrix has been computed from Eq. (3), it can be used to predict the evolution of dynamical variables by
$\boldsymbol{u}_{p}(t_{i+1})=[\boldsymbol{r}(t_{i+1})]^{{\textnormal{T}}}\,\boldsymbol{W}_{out}.$ (4)
@@
In the closed-loop configuration, we recursively predict the system’s dynamics from Eq. (4) without additional samples from a training data set. Given an initial condition, this allows for an autonomous prediction on the unseen data set. Because reservoir computers have a symmetry (Eq. (1)), we add a constant output bias of 1 to the reservoir activation states to break the inherent symmetry of the reservoir architecture Huhn and Magri (2020b); Herteux and Räth (2020), effectively replacing $\boldsymbol{r}\rightarrow[\boldsymbol{r},1]$. A summary of the reservoir computing procedure is shown in Fig. 3, and compared with its quantum counterpart, which will be introduced in Sec. II.3. The performance of reservoir computers critically depends on the set of chosen hyper-parameters. We, therefore, use grid search and Bayesian optimization to tune the hyperparameters Snoek et al. (2012). We have used $scikit-learn$ library Pedregosa et al. (2011) and recycle validation techniques Racca and Magri (2021) for hyperparameter tuning. The hyperparameters and performance metrics for different chaotic systems are reported in Sec. IV.

### v2

# Prediction of chaotic dynamics and extreme events: A recurrence-free quantum reservoir computing approach

@@ Abstract
In chaotic dynamical systems, extreme events manifest in time series as unpredictable large-amplitude peaks. Although deterministic, extreme events appear seemingly randomly, which makes their forecasting difficult. By learning the dynamics from observables (data), reservoir computers can time-accurately predict extreme events and chaotic dynamics, but they may require many degrees of freedom (large reservoirs). In this paper, by exploiting quantum-computer ansätze and entanglement, we design reservoir computers with compact reservoirs and accurate prediction capabilities. First, we propose the recurrence-free quantum reservoir computer (RF-QRC) architecture. By developing ad hoc quantum feature maps and removing recurrent connections, the RF-QRC has quantum circuits with smaller depths. This allows the RF-QRC to scale well with higher-dimensional chaotic systems, which makes it suitable for hardware implementation. Second, we forecast the temporal chaotic dynamics and their long-term statistics of low- and higher-dimensional dynamical systems. We find that RF-QRC requires smaller reservoirs than classical reservoir computers for higher-dimensional systems and the same predictability. Third, we apply the RF-QRC to the time prediction of extreme events in a model of a turbulent shear flow with turbulent bursts. We find that the RF-QRC has longer predictability than the classical reservoir computer for extreme events forecasting. The results and analyses indicate that quantum-computer ansätze offers nonlinear expressivity and computational scalability, which are useful for forecasting chaotic dynamics and extreme events. This work opens new opportunities for using quantum machine learning on near-term quantum computers.
@@ I Introduction:
Data-driven prediction of chaotic dynamical systems has gained significant interest in the last decade [1, 2]. A data-driven model learns the solution of a dynamical system from data with the goal of predicting its temporal evolution. Weather and climate predictions [3, 4], financial time-series forecasting [5], thermoacoustics [6], turbulence [7, 8], among many others, are examples of chaotic dynamical systems. The use of data-driven methods for predicting nonlinear systems has been motivated mainly by two reasons: (a) the availability of large amounts of data for these systems, and (b) the difficulty of predicting temporal dynamics due to chaos. In chaotic systems, a small change in the initial conditions can drastically change the dynamical system’s solution.
@@
Because chaotic dynamics is a sequential-data problem, Recurrent Neural Networks (RNNs) are a natural choice for chaotic forecasting. RNNs are machine learning models that introduce recurrence in conventional neural network architectures. RNNs are employed in time-series forecasting, modeling nonlinear dynamical systems and chaos, and extreme event predictions [9, 7, 10, 11, 12, 13]. On the one hand, the computational capabilities of RNNs are excellent. On the other hand, RNNs are difficult to train as they require backpropagation through time at each time step [14]. Because RNNs are designed to learn correlations using internal hidden states and long-lasting time dependencies, the training through backpropagation can be difficult for long time-series forecasting [9].
@@
Reservoir Computing (RC) bypasses the problem of backpropagation by introducing a static recurrence—the reservoir—and training with ridge regression. This avoids backpropagation at each time step. Reservoir computing is a unified computing framework that was introduced as ’Echo State Networks’ [15] and ’Liquid State Machines’ [16]. In this study, we work with Echo State Networks, which are a type of reservoir computing approach, because of their potential in forecasting chaotic dynamics [11, 9, 17], extreme events predictions [8, 18, 19] and the stability properties [12, 6]. (For brevity, we will refer to echo state networks as reservoir computing in this work). Despite their excellent forecasting abilities, the prediction capabilities of reservoir computers are sensitive to the choice of hyperparameters and the reservoir size, which is constrained by the computational resources and memory of current classical computers.
@@
Improving the performance of reservoir computing typically requires higher computational resources. Quantum computers hold the potential to provide exponential speed-up over classical computers for certain computational tasks [20]. This speed-up is also known as ’quantum advantage’. There are already algorithms that have theoretically proven this concept. For example, the prime factorization problem, which is the basis of most modern encryption and is NP-hard classically, can be solved in polynomial time using quantum computing principles [21]. The second example is the matrix inversion problem for solving linear systems of equations, which is a key step in many numerical solvers. Proposed quantum algorithms for matrix inversion have the potential to provide quadratic speed-up over conventional classical methods [22, 23]. The underlying phenomena that enable the quantum advantage are entanglement, superposition, and interference effects in quantum computers. Recently, quantum computing has been applied in the area of machine learning. The current quantum hardware, however, is limited by the maximum number of qubits in the order of $\sim$ $10^{2}-10^{3}$ with environmental noise and decoherence. In the current Noisy Intermediate Scale Quantum (NISQ)-era [24], hybrid quantum-classical methods are one of the leading candidates with prospects of achieving a quantum advantage by employing methods such as Variational Quantum Algorithms and Quantum Circuit Learning [25, 26]. The hybrid quantum-classical methods that can be used for the prediction of chaotic dynamics and extreme events are inspired by classical machine learning architectures such as RNNs.
@@
Quantum Reservoir Computing (QRC) is a type of Quantum Machine Learning model that combines both frameworks, including Quantum Computing and Reservoir Computing. The idea of QRC is to use quantum dynamics to enhance the reservoir implementation [27, 28, 29]. Some numerical implementations of QRC have shown that the quantum systems of 5-7 qubits have comparable computational capabilities to the classical reservoir sizes of 100-500 [27]. Previous works have proposed hybrid QRC architectures for time series predictions and chaos modeling [30, 31, 32, 33]. In hybrid QRC, the input data is encoded in the form of Bloch-sphere rotation angles in individual qubits along with reservoir states, which are the probability amplitudes. The resulting quantum states undergo a unitary operation at each time step [30]. In another approach, a reservoir state is associated with the density operator of the encoded quantum state, and reservoir states correspond to measured expectation values for each qubit [32, 34]. Although it has been conjectured that quantum reservoir computers are potential candidates for providing a quantum advantage in the near-term NISQ era, only a few studies have been conducted on analyzing higher-dimensional chaotic systems [35], and long-term statistical predictions using QRC. The performance of classical reservoir computers is often limited by classical computers’ memory [9], which constrains the predictability of these models. Reservoir computers require a high dimensional feature space, which increases with the nonlinearity and the dimensions of the physical system. Consequently, using QRC to predict high dimensional nonlinear dynamical systems that require a large reservoir size is key to realizing a quantum advantage of QRC in predicting chaos.
@@
In this work, we build up on the concept of hybrid QRC that encodes reservoir states as probability amplitudes of the quantum system [30]. We use gate-based quantum computing to implement this architecture and use it for high-dimensional chaotic time-series forecasting and extreme-event predictions. More specifically, we investigate the three-dimensional Lorenz-63, ten-dimensional Lorenz-96 systems [36] and the MFE (Moehlis, Faisst, and Eckhardt) [37] model for time-accurate, statistical predictions, and extreme events forecasting. Extreme events are sudden and unmitigated changes in the observables of chaotic flows, while long-term statistical prediction is also an important metric in forecasting chaos.
@@
We benchmark our classical and quantum reservoir networks by studying them for low-order models and then extend our analysis to incorporate higher-dimensional dynamics. Both classical and quantum networks are used for extreme event predictions in a reduced order model of a shear flow between infinite plates subjected to a sinusoidal body forcing, also referred to as ’Moehlis, Faisst, and Eckhardt (MFE)’ [37]. We assess the performance using multiple performance metrics: first, by considering short-term time series prediction capabilities that are quantified using the Predictability Horizon (PH) and Valid Prediction Time (VPT) [9, 38], and, second, by comparison of long-term statistics that are quantified using statistical measures such as Probability Density Function (PDF) and F-Score [12, 13].
@@
This paper is structured as follows. In Sec. II, we present the classical and proposed quantum reservoir architectures and their comparison. Different quantum reservoir architectures refer to the different ansätze chosen for forming a reservoir. Second, we apply both classical and quantum reservoir architectures on three-dimensional Lorenz-63 and ten-dimensional Lorenz-96 systems. We increase the dimensionality and complexity of the dynamical systems progressively for the analysis. Third, we compare the results of the best quantum feature map with the best classical reservoir applied to the MFE model to analyze the extreme-event prediction capabilities of both architectures. The numerical models of different dynamical systems and the results for the reservoir predictions are presented in Sec. IV. Finally, in Sec. V, we conclude our findings by highlighting the potential of QRC in chaotic time series forecasting along with its associated challenges and the direction of future works.
@@ II.1 Classical Reservoir Computing (CRC)
Unlike other RNNs, reservoir computing is a type of RNN framework that does not require backpropagation through time, which represents a major computational advantage [15, 13]. As a result, the training cost is substantially reduced and the objective function is minimized through a simple linear ridge regression.
Figure 1: Schematic representation of a classical reservoir computer [15]. The input data $\pmb{u}_{in}$ is mapped to the reservoir matrix via $\pmb{W}_{in}$. The reservoir neuron connections governed by $\pmb{W}$ matrix allow the flow of information between neurons. The linear readout layer using the trained $\pmb{W}_{out}$ matrix is used to make output predictions $\pmb{u}_{p}$.
Figure 2: Schematic representation of reservoir computers. Open-loop training phase.
Figure 3: Schematic representation of reservoir computers. Closed-loop prediction phase.
@@
Let $\{\bm{\hat{u}}(t_{0}),\bm{\hat{u}}(t_{1}),\ldots,\bm{\hat{u}}(t_{N_{tr}})\}$ be a training data set of a dynamical state $\hat{\pmb{u}}\,\epsilon\,\,\mathbb{R}^{{N}_{u}}$ , which is known at $N_{tr}+1$ steps in time. Generally, it is recommended to choose a set of variables that is (re-)scaled by the range component-wise, which is indicated by ( $\hat{}$ ) on the training data set [39]. In reservoir computing, these states are mapped onto higher-dimensional reservoir states of an $N_{r}$-dimensional vector space. For a randomly chosen initial reservoir state $\bm{r}(t_{0})$, we recursively compute a set of $N_{tr}$ reservoir vectors $\pmb{r}(t_{1}),\ldots,\pmb{r}(t_{N_{tr}})$ as
$\pmb{\hat{r}}(t_{i+1})$ $=\tanh\,(\pmb{W}_{in}\pmb{\hat{u}}_{in}(t_{i})+\pmb{W}\pmb{r}(t_{i})),$ (1)
$\pmb{r}(t_{i+1})$ $=(1-\varepsilon)\,\pmb{r}(t_{i})+\varepsilon\,\pmb{\hat{r}}(t_{i+1}).$ (2)
@@
Here, the state $\hat{\pmb{r}}\,\epsilon\,\,\mathbb{R}^{{N}_{r}}$ represent a state-vector containing reservoir activation states, $\pmb{W}_{in}\,\epsilon\,\,\mathbb{R}^{{N}_{r}\,\times{N}_{u}}$ is the input matrix and $\pmb{W}\,\epsilon\,\,\mathbb{R}^{{N}_{r}\,\times{N}_{r}}$ is the reservoir weight matrix [15]. The $\tanh$ function is applied component-wise to provide a nonlinear activation. The matrices $\pmb{W}_{in}$ and $\pmb{W}$ are pseudo-randomly generated and constant throughout the training and prediction phase. The elements of $\pmb{W}_{in}$ are sampled from a uniform distribution in $[-{\sigma_{in}},{\sigma_{in}}]$, where ${\sigma_{in}}$ is the input-scaling. The reservoir weight matrix $\pmb{W}$ is an Erdös-Rényi matrix (Erdös-Rényi is a model for generating random graphs or the evolution of a random network in the field of graph theory), with average connectivity between each neuron varied by input reservoir density $D$. The spectral radius of the reservoir state matrix $\rho$ is given by its maximal eigenvalue. The matrix $\bm{W}$ is rescaled such that $\rho\leq 1$ [39, 15]. Effectively, this guarantees a rapid decay of temporal correlations between successive reservoir states and helps to avoid overfitting. Whereas Eq. (1) is the nonlinear activation step, Eq. (2) combines the linear memory of the reservoir with the nonlinear activation, as parametrized by the leaking rate $\varepsilon$. This additional post-processing step is known as memory non-linearity tradeoff [39].
[…]
@@
where $\beta$ is a user-defined Tikhonov regularization parameter to prevent overfitting, $\pmb{I}$ is the identity matrix, and $\pmb{{U}}^{{\textnormal{T}}}_{d}\,\,\epsilon\,\,\mathbb{R}^{{N}_{u}\,\times{N}_{tr}}$ is the horizontal concatenation of the output data. Once the output matrix has been computed from Eq. (3), it can be used to predict the evolution of dynamical variables by
$\pmb{u}_{p}(t_{i+1})=[\pmb{r}(t_{i+1})]^{{\textnormal{T}}}\,\pmb{W}_{out}.$ (4)
@@
In the closed-loop configuration, we recursively predict the system’s dynamics from Eq. (4) without additional samples from a training data set. Given an initial condition, this allows for an autonomous prediction on the unseen data set. Because reservoir computers have a symmetry (Eq. (1)), we add a constant output bias of 1 to the reservoir activation states to break the inherent symmetry of the reservoir architecture [40, 41], effectively replacing $\bm{r}\rightarrow[\bm{r},1]$. A summary of the reservoir computing procedure is shown in Fig. 4, and compared with its quantum counterpart, which will be introduced in Sec. II.3. The performance of reservoir computers critically depends on the set of chosen hyper-parameters. We, therefore, use grid search and Bayesian optimization to tune the hyperparameters [42]. We have used the $scikit-learn$ library [43] and recycle validation techniques [13] for hyperparameter tuning. The hyperparameters and performance metrics for different chaotic systems are reported in Sec. IV.